OpenStudy (anonymous):

In the numerical sentence below, A = 8, D = , and F = . What are the values B, C, and E if the sentence below is equivalent to the numerical sentence, 2 × 4 = 8? B = C = E =

OpenStudy (anonymous):

D= 1/2 F= 1/4